40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 04:08:53 2015, Jenne, Update, LSC, New Locking Paradigm? 
    Reply  Wed Feb 11 22:13:44 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes 
       Reply  Thu Feb 12 01:43:09 2015, rana, Update, LSC, New Locking Paradigm - LSC model changes TF.gif
       Reply  Thu Feb 12 11:14:29 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra ALS_REFL_comboLockingCartoon_11Feb2015.PDFALS_REFL_comboLocking_11Feb2015.zip
          Reply  Thu Feb 12 11:59:58 2015, Koji, Update, LSC, New Locking Paradigm - Loop-gebra 
             Reply  Thu Feb 12 19:18:49 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra 
                Reply  Fri Feb 13 03:28:34 2015, rana, Update, LSC, New Locking Paradigm - Loop-gebra 
                   Reply  Sat Feb 14 20:20:24 2015, Jenne, Update, LSC, ALS fool cartoon ALSfool_LoopDiagram.pngALSfool_LoopDiagram.graffle.zip
                      Reply  Tue Feb 17 04:04:32 2015, Jenne, Update, LSC, ALS fool math OpenLoopGainComparison_16Feb2015.pngCancellationTFs_DecouplingIsArefl.png
                         Reply  Tue Feb 17 16:36:08 2015, Jenne, Update, LSC, ALS fool math ALSfool_LoopDiagram_17Feb2015.pngALS_REFL_comboLocking_16Feb2015.nb.zip
       Reply  Thu Feb 12 22:28:16 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes, screens modified 
          Reply  Sat Mar 7 19:15:17 2015, Jenne, Update, LSC, Modified zero crossing triggering LatchLogic_6March2015.png
    Reply  Thu Feb 12 03:43:54 2015, ericq, Update, LSC, 3F PRMI at zero ALS CARM Feb12_negativeTR.png
Message ID: 11001     Entry time: Wed Feb 11 04:08:53 2015     Reply to this: 11007   11010
Author: Jenne 
Type: Update 
Category: LSC 
Subject: New Locking Paradigm? 

[Rana, Jenne]

While meditating over what to do about the fact that we can't seem to hold PRMI lock while reducing the CARM offset, we have started to nucleate a different idea for locking

We aren't sure if perhaps there is some obvious flaw (other than it may be tricky to implement) that we're not thinking about, so we invite comments.  I'll make a cartoon and post it tomorrow, but the idea goes like this.....

Can we use ALS to hold both CARM and DARM by actuating on the ETMs, and sit at (nominally) zero offset for all degrees of freedom?  PRMI would need to be stably held with 3f signals throughout this process. 

1) Once we're close to zero offset, we should see some PDH signal in REFL11.  With appropriate triggering (REFLDC goes low, and REFL11I crosses zero), catch the zero crossing of REFL11I, and feed it back to MC2. We may want to use REFL11 normalized by the sum of the arm transmissions to some power (1, 0.5, or somewhere in between may maximize the linear range even more, according to Kiwamu).  The idea (very similar to the philosophy of CESAR) is that we're using ALS to start the stabilization, so that we can catch the REFL11 zero crossing. 

2) Now, the problem with doing the above is that actuating on the mode cleaner length will change the laser frequency.  But, we know how much we are actuating, so we can feed forward the control signal from the REFL11 carm loop to the ALS carm loop.  The goal is to change the laser frequency to lock it to the arms, without affecting the ALS lock.  This is the part where we assume we might be sleepy, and missing out on some obvious reason why this won't work.

3) Once we have CARM doubly locked (ALS pushing on ETMs, REFL11 pushing on MC/laser frequency), we can turn off the ALS system. Once we have the linear REFL11 error signal, we know that we have enough digital gain and bandwidth to hold CARM locked, and we should be able to eek out a slightly higher UGF since there won't be as many digital hops for the error signal to transverse. 

4) The next step is to turn on the high bandwidth common mode servo.  If ALS is still on at this point, it will get drowned out by the high gain CM servo, so it will be effectively off. 

5) Somewhere in here we need to transition DARM to AS55Q.  Probably that can happen after we've turned on the digital REFL11 path, but it can also probably wait until after the CM board is on.

The potential show-stoppers:

Are we double counting frequency cancellation or something somewhere?  Is it actually possible to change the laser frequency without affecting the ALS system?

Can we hold PRMI lock on 3f even at zero CARM offset?  Anecdotally from a few trials in the last hour or so, it seems like coming in from negative carm offset is more successful - we get to slightly higher arm powers before the PRMI loses lock.  We should check if we think this will work in principle and we're just saturating something somewhere, or if 3f can't hold us to zero carm offset no matter what.

A note on technique:  We should be able to get the transfer function between MC2 actuation and ALS frequency by either a direct measurement, or Wiener filtering.  We need this in order to get the frequency subtraction to work in the correct units.

ELOG V3.1.3-